OpenStudy (anonymous):

2x - y = 3 4x = 6 + 2y

OpenStudy (zzr0ck3r):

they are the same -2(2x-y=3) = -4x+2y = -6 and -1(4x= 6+2y) = -4x +2y = -6 so they are equal for all points im not sure if that is what you are asking, but im bored so I tried to guess:)